# Geodesic loops and orthogonal geodesic chords without self-intersections

### v1

@@ 3. Geodesics with product boundary conditions
The following concept is introduced in [2, Sec.4], cf. also [10]. Let $M$ be a differentiable manifold. Let $\mathcal{P}\subset M\times M$ be a compact submanifold. This is called a general boundary condition on $M$ if for some Riemannian metric $g$ on $M$ the pseudo-Riemannian metric $\overline{g}=g\oplus(-g)$ on $M\times M$ induces a non-degenerate pseudo-Riemannian metric on $\mathcal{P},$ cf.  [2, Def.4.2]. It is shown in [2, Prop.4.1] that the subset $\mathcal{A}_{\mathcal{P}}\subset\mathcal{G}(M)$ of the set of Riemannian metrics on $M$ for which the restriction of the pseudo Riemannian metric $\overline{g}$ onto the submanifold $\mathcal{P}$ is non-degenerate, is an open subset. The second fundamental form of $\mathcal{P}$ in the normal direction $\eta\in T\mathcal{P}^{\perp}$ is given by
@@
cf. [2, p.346]. Here $\overline{\nabla}$ is the Levi-Civita connection on $M\times M$ with metric $\overline{g}.$ The vector field $\overline{w}$ is a smooth extension of $w$ to a smooth vector field along $\mathcal{P}.$ Since the restriction of $\overline{g}$ to $\mathcal{P}$ is non-degenerate we can identify the symmetric bilinear form with the self-adjoint shape operator $(S_{\eta}^{\mathcal{P}})_{p}:T_{p}\mathcal{P}\longrightarrow T_{p}\mathcal{P}$ with $\overline{g}((S_{\eta}^{\mathcal{P}})_{p}v,w)=S_{\eta}^{\mathcal{P}}(v,w)$ for all $v,w\in T_{p}\mathcal{P}.$ Let $H^{1}(I,M)$ be the set of all curves of Sobolev class $H^{1}$ in $M,$ i.e. of absolutely continuous curves with finite energy. Then the subspace
@@
[…]
A geodesic $\gamma:I\longrightarrow M$ is called a $(g,\mathcal{P})-$geodesic if $(\gamma^{\prime}(0),\gamma^{\prime}(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$ Here the orthogonality is with respect to the metric $\overline{g}.$ It is shown in [2, Sec. 5] that a curve $\gamma$ is a critical point of the energy functional $E_{g}:\Omega_{\mathcal{P}}(M)\longrightarrow\mathbb{R},$
[…]
@@
A general boundary condition is called admissible if for every $g_{0}\in\mathcal{A}_{\mathcal{P}}$ there exists an open neighborhood $\mathcal{V}$ of $g_{0}$ in $\mathcal{A}_{\mathcal{P}}$ and a positive real number $a>0$ such that for all $g\in\mathcal{V}$ and all non-trivial $(g,\mathcal{P})$-geodesics $\gamma$ we have: $L_{g}(\gamma)\geq a,$ cf. [2, Def.4.5]. Here a non-trivial $(g,\mathcal{P})$ geodesic is a non-constant curve.
@@
In particular a $(g,N_{0}\times N_{1})$-geodesic is a geodesic $c:[0,1]\longrightarrow M$ with $c(l)\in N_{l},c^{\prime}(l)\in T_{c(l)}^{\perp}N_{l},l=0,1.$ The product metric $g\oplus(-g)$ on $N_{0}\times N_{1}$ is non-degenerate, hence a product boundary condition is always non-degenerate, as defined in [2, Sec.4]. In case $\mathcal{P}=N\times N$ the $(g,N\times N)$-Jacobi fields are Jacobi fields $Y\in T_{\gamma}\Omega_{N\times N}$ with
@@
(a) Let $(M,g)$ be a compact Riemannian manifold (without boundary) of dimension $n\geq 2.$ If $p,q\in M$ are two points, where we also allow $p=q,$ then $\{p\}\times\{q\}$ is an admissible product boundary condition.
(b) If $(M,g)$ is a compact Riemannian manifold of dimension $n\geq 2$ with smooth boundary $\partial M$ then $\partial M\times\partial M$ is an admissible product boundary condition.
(c) If $(M,g)$ is a compact Riemannian manifold (without boundary) of dimension $n\geq 2$ and if $N\subset M$ is a compact submanifold of codimension one then $N\times N$ is an admissible product boundary condition.
@@
(b) and (c). Here the normal exponential map of the submanifold $\partial M$ resp. $N$ is defined $\exp^{\perp}:\nu(\partial M)\longrightarrow M,$ resp. $\exp^{\perp}:\nu(N)\longrightarrow M.$ Here $\nu(\partial M)$ is the subset of the normal bundle of $\partial M$ of normal vectors pointing inward and $\nu(N)$ is the normal bundle of $N\subset M.$ Then there is a positive and continuous function $g\in\mathcal{G}^{r}(M)\mapsto\mathrm{inj}^{\perp}(g)$ such that the restriction onto the normal vectors of length $\leq\mathrm{inj}^{\perp}$ is a diffeomorphism onto its image. Any non-trivial $(g,\partial M\times\partial M)$-geodesic resp. $(g,N\times N)$-geodesic $\gamma:[0,1]\longrightarrow M$ satisfies $L_{g}(\gamma)\geq 2\mathrm{inj}^{\perp}(g).$
@@
It is a standard construction that one can add a cylindrical neck to the manifold $M$ with smooth boundary. I.e. there exists a Riemannian manifold $M_{0}$ with metric $g_{0}$ such that $M$ is an open submanifold and $g$ is the restriction of $g_{0}$ to $M.$ And for some positive $\epsilon>0$ the normal exponential map of the hypersurface $\partial M$ defines an injective immersion $\phi:(t,x)\in[-\epsilon,\epsilon]\times\partial M\longmapsto M_{0}.$ The Riemannian metric then has the form $dt^{2}+h_{t}$ for a smooth family $t\in[-\epsilon,\epsilon]\mapsto h_{t}\in\mathcal{G}(\partial M)$ of Riemannian metrics $h_{t}$ on $\partial M$ and $h_{t}=h_{1}$ for all $t\in[2/3,1].$ Hence the set $\phi([2/3,1]\times\partial M)$ is the cylindrical end, which is isometric to the Riemannian product $[2/3,1]\times(\partial M,h_{1}).$ This construction can be found for example in [8, Section 5].
Then one can take two copies of $(M_{0},g_{0})$ identified along the boundary $\partial M_{0},$ hence we obtain a compact Riemannian manifold $(M_{1},g_{1})$ without boundary such that $\partial M_{0}$ is a totally geodesic submanifold. ∎
@@
In the setting of Lemma 5 let $\phi:\mathcal{G}(M_{1})\longrightarrow\mathcal{G}(M)$ be the map defined by restricting the Riemannian metric $g$ on $M_{1}$ to the submanifold $M,$ i.e. $\phi(g)=g|M.$ The following statements are an immediate consequence of the definition of the strong topology, in particular of the definition of a strong basic neighborhood cf. [12, Section 2.1] applied to the space of Riemannian metrics on $M$ resp. $M_{1}.$ After fixing a Riemannian metric the space $\mathcal{G}(M)$ of Riemannian metrics of class $C^{r},r\geq 2$ on the manifold $M$ endowed with with the strong topology can be seen as an open subset in the Banach space of $C^{r}$-sections of the bundle of symmetric bilinear forms on the manifold $M.$ If $\mathcal{U}$ is an open and dense subset of $\mathcal{G}(M_{1})$ then $\phi(\mathcal{U})$ is an open and dense subset of $\mathcal{G}(M).$ And if $\mathcal{V}$ is a residual subset of $\mathcal{G}(M_{1}),$ then the restriction $\phi(\mathcal{V})$ is a residual subset of $\mathcal{G}(M).$
@@
(b) Let $(M,g)$ a compact Riemannian manifold of dimension $n\geq 2$ and let $N\subset M$ be a compact submanifold of codimension $1$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics on $M$ such that for any $g\in\mathcal{G}^{*}[L]$ no $(g,N\times N)$-geodesic of length $\leq L$ is part of a closed geodesic.
@@
For $L>0$ let $\mathcal{G}^{*}[L]$ be the set of Riemannian metrics for which all closed geodesics with energy $\leq L$ are non-degenerate and simple, cf. [18, Thm.1] such that no restriction onto a subinterval defines - up to linear reparametrization - a geodesic segment between $p$ and $q$ in case (a) resp. a $(g,N\times N)$-geodesic in case (b). This set is open, we have to show that it is a dense subset of $\mathcal{G}(M).$ A metric $g\in\mathcal{G}^{*}[L]$ has only finitely many geometrically distinct prime closed geodesics $c_{j}:S^{1}\longrightarrow M,j=1\ldots,k$ of energy $\leq L$ and there is an open subset $\mathcal{U}\subset\mathcal{G}^{*}[L]$ and continuous functions $g\in\mathcal{U}\mapsto v_{j}(g)\in TM$ such that the curve $t\in\mathbb{R}\mapsto c_{j}(t)=\Phi_{g}^{t}(v_{j}(g)),t\in S^{1}$ is a closed geodesic of $g$ and there are - up to geometric equivalence - no further closed geodesics of length $\leq L$ of the metric $g,$ cf. [1, p.12-13], resp. the Proof of Part (b) of Theorem 4. Here $\Phi^{t}_{g}:TM\longrightarrow TM,t\in\mathbb{R}$ denotes the geodesic flow, i.e. for $v\in T_{p}M,t\in\mathbb{R}\mapsto\Phi_{g}^{t}(v)\in TM$ is the periodic orbit in the tangent bundle whose projection on the manifold is the geodesic $c:\mathbb{R}\longrightarrow M$ with initial point $p=c(0)$ and initial direction $v=c^{\prime}(0).$
@@
Let $D^{n-1}_{2\eta}=\{x\in\mathbb{R}^{n-1}\,;\,\|x\|\leq 2\eta\}$ be the Euclidean $(n-1)$-disc of radius $2\eta$ for $\eta>0.$ On $(t,x)=(t,x_{1},\ldots,x_{n-1})\in[-2\eta,2\eta]\times D^{n-1}_{2\eta}$ a Riemannian metric is given by
@@
$\frac{1}{(1-\epsilon f^{\prime}(t))^{2}}\left(ds^{2}+2\epsilon f(t)dsdx_{1}+\epsilon^{2}f^{2}(t)dx_{1}^{2}\right)+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}\,.$
@@
is sufficently close go $g$ in the $C^{r}$-topology for $r\geq 2$ and the the curve $x_{1}=\ldots=x_{n-1}=0$ coincides with $\gamma$ and is a geodesic of $\overline{g}$ parametrized by arc length. At the point $p=\gamma(0)$ the curve is orthogonal to the hypersurface $s=0$ but not to the hypersurface $\Sigma=\{t=0\}.$
@@
(c) Let $M$ be a compact differentiable manifold (without boundary) of dimension $n\geq 3$ and let $N\subset M$ be a compact submanifold of codimension $1$ Then for a $C^{k}$-generic metric on $M$ with $k\geq 2$ all $(g,N\times N)$-geodesics are non-degenerate and simple, and any two geometrically distinct $(g,N\times N)$-geodesics do not intersect.
@@
It follows from Lemma 6 and from the result [2, Thm.5.1], cf. also [2, p.347], that for $L>0$ the set $\mathcal{G}_{\mathcal{P}}(M)[L]$ of metrics $g\in\mathcal{G}_{\mathcal{P}}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and are not part of a closed geodesic is an open and dense subset of $\mathcal{G}_{\mathcal{P}}=\mathcal{G}_{\mathcal{P}}(M).$ This holds since in all three cases (a),(b),(c) the boundary conditions $\mathcal{P}$ are non-degenerate and admissible, cf. Lemma 4.
Here one has to note that in case (b) one has to consider the differentiable manifold $M$ to be a subset of a compact manifold $M_{1}$ without boundary, cf. Lemma 5. Then one can apply [2, Thm.5.1] and conclude that the set $\mathcal{G}_{\partial M\times\partial M}(M_{1})[L]$ of metrics $g\in\mathcal{G}(M_{1})$ for which all $(g,\partial M\times\partial M)$-geodesics of length $\leq L$ are non-degenerate is an open and dense subset of $\mathcal{G}_{\partial M\times\partial M}(M_{1}).$ Then the set $\mathcal{G}_{\partial M\times\partial M}(M)$ contains all restrictions $g|M$ of metrics $g\in\mathcal{G}_{\partial M\times\partial M}(M_{2})[L],$ and is therefore also dense in $\mathcal{G}(M),$ cf. Remark 3.
@@
We conclude from [10, Thm.2.4] that the energy functional $E_{g}:c\in\Omega_{\mathcal{P}}(M)\longmapsto E_{g}(c)\in\mathbb{R}$ satisfies the Palais-Smale condition. We can apply this result since $M$ is compact, hence any Riemannian metric on $M$ is complete.
@@
The self-adjoint operator $H_{c}$ representing the hessian $d^{2}E_{g}(c)$ at a $(g,\mathcal{P})$-geodesic $c$ is of the form $H_{c}=\mathbbm{1}+k_{c},$ i.e. the sum of the identity and a compact operator $k_{c}.$ Since 00 is not an eigenvalue the inverse $H_{c}^{-1}$ is a bounded linear operator. Therefore we can use the implicit function theorem for Banach manifolds, cf. [5, Thm.15.1] and conclude: There is a neigbhourhood $\mathcal{U}$ of $g$ and continuous functions $g\in\mathcal{U}\longmapsto c_{j}[g]\in\Omega_{\mathcal{P}}(M),j=1,\ldots,k$ such that $c_{j}[g]$ is a $(g,\mathcal{P})$-geodesic of $g\in\mathcal{U}.$
One can choose the open neighborhhod $\mathcal{U}$ sufficiently small such that there are no further $(g,\mathcal{P})$-geodesics of $g$ with $E_{g}(\gamma)\leq L,$ cf. the argument in [1, p.12-13].
@@
We first assume $p=q.$ For $L>0$ for a Riemannian metric $g\in\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ there are only finitely many geometrically distinct and non-degenerate geodesic loops $\gamma_{1},\ldots,\gamma_{N}$ of length $\leq L.$ Assume that - after a possible renumbering - the geodesic loops $\gamma_{1},\ldots,\gamma_{m},m\leq N$ are the prime geodesic loops of length $\leq L.$ Here a geodesic loop $\gamma:[0,1]/\{0,1\}\longrightarrow M$ is prime, if there is no $t_{1}\in(0,1)$ with $p=c(t_{1}).$ Let $v_{j,l}=\gamma_{j}^{\prime}(l),j=1,\ldots,m,l\in\{0,1\}$ be the sequence of tangent vectors of the prime geodesic loops at $p.$ We use the perturbation result Lemma 3 which shows that there is an open and dense subset $\mathcal{G}_{p}[L]\subset\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for which all geodesic loops at $p$ of length $\leq L$ are prime. Note here that if $m<N,$ then there the geodesic loops $\gamma_{m+1},\ldots,\gamma_{N}$ are not prime, but their prime parts are loops $\gamma_{j},1\leq j\leq m.$ Since after the perturbation these prime geodesic loops cannot be concatenated to produce a non-prime geodesic loop the perturbed loops $\gamma_{j}[g],j\geq m+1$ have to be prime, too.

### v2

@@ 3. Geodesics with product boundary conditions
The following concept is introduced in [4, Sec.4], cf. also [12]. Let $M$ be a differentiable manifold. Let $\mathcal{P}\subset M\times M$ be a compact submanifold. This is called a general boundary condition on $M$ if for some Riemannian metric $g$ on $M$ the pseudo-Riemannian metric $\overline{g}=g\oplus(-g)$ on $M\times M$ induces a non-degenerate pseudo-Riemannian metric on $\mathcal{P},$ cf.  [4, Def.4.2]. It is shown in [4, Prop.4.1] that the subset $\mathcal{A}_{\mathcal{P}}\subset\mathcal{G}(M)$ of the set of Riemannian metrics on $M$ for which the restriction of the pseudo Riemannian metric $\overline{g}$ onto the submanifold $\mathcal{P}$ is non-degenerate, is an open subset. The second fundamental form of $\mathcal{P}$ in the normal direction $\eta\in T\mathcal{P}^{\perp}$ is given by
@@
cf. [4, p.346]. Here $\overline{\nabla}$ is the Levi-Civita connection on $M\times M$ with metric $\overline{g}.$ The vector field $\overline{w}$ is a smooth extension of $w$ to a smooth vector field along $\mathcal{P}.$ Since the restriction of $\overline{g}$ to $\mathcal{P}$ is non-degenerate we can identify the symmetric bilinear form with the self-adjoint shape operator $(S_{\eta}^{\mathcal{P}})_{p}:T_{p}\mathcal{P}\longrightarrow T_{p}\mathcal{P}$ with $\overline{g}((S_{\eta}^{\mathcal{P}})_{p}v,w)=S_{\eta}^{\mathcal{P}}(v,w)$ for all $v,w\in T_{p}\mathcal{P}.$ Let $H^{1}(I,M)$ be the set of all curves of Sobolev class $H^{1}$ in $M,$ i.e. of absolutely continuous curves with finite energy. Then the subspace
@@
$\langle v,w\rangle_{1}=\int_{0}^{1}g\left(v(t),w(t)\right)\,dt+\int_{0}^{1}g\left(\nabla v(t),\nabla w(t)\right)\,dt\,.$
A geodesic $\gamma:I\longrightarrow M$ is called a $(g,\mathcal{P})$-geodesic if $(\gamma^{\prime}(0),\gamma^{\prime}(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$ Here the orthogonality is with respect to the metric $\overline{g}.$ It is shown in [4, Sec. 5] that a curve $\gamma$ is a critical point of the energy functional $E_{g}:\Omega_{\mathcal{P}}(M)\longrightarrow\mathbb{R},$
(1) $E_{g}(\gamma)=\frac{1}{2}\int_{0}^{1}g(\gamma^{\prime}(t),\gamma^{\prime}(t))\,dt$
@@
A general boundary condition is called admissible if for every $g_{0}\in\mathcal{A}_{\mathcal{P}}$ there exists an open neighborhood $\mathcal{V}$ of $g_{0}$ in $\mathcal{A}_{\mathcal{P}}$ and a positive real number $a>0$ such that for all $g\in\mathcal{V}$ and all non-trivial $(g,\mathcal{P})$-geodesics $\gamma$ we have: $L_{g}(\gamma)\geq a,$ cf. [4, Def.4.5]. Here a non-trivial $(g,\mathcal{P})$ geodesic is a non-constant curve.
@@
In particular a $(g,N_{0}\times N_{1})$-geodesic is a geodesic $c:[0,1]\longrightarrow M$ with $c(l)\in N_{l},c^{\prime}(l)\in T_{c(l)}^{\perp}N_{l},l=0,1.$ The product metric $g\oplus(-g)$ on $N_{0}\times N_{1}$ is non-degenerate, hence a product boundary condition is always non-degenerate, as defined in [4, Sec.4]. In case $\mathcal{P}=N\times N$ the $(g,N\times N)$-Jacobi fields are Jacobi fields $Y\in T_{\gamma}\Omega_{N\times N}$ with
@@
The following sets $\mathcal{P}$ are admissible product boundary conditions:
(a) Let $(M,g)$ be a compact Riemannian manifold (without boundary) of dimension $n\geq 2.$ If $p,q\in M$ are two points, where we also allow $p=q,$ then $\mathcal{P}=\{p\}\times\{q\}$ is an admissible product boundary condition.
(b) If $(M,g)$ is a compact Riemannian manifold of dimension $n\geq 2$ with smooth boundary $\partial M,$ then $\mathcal{P}=\partial M\times\partial M$ is an admissible product boundary condition.
(c) If $(M,g)$ is a compact Riemannian manifold (without boundary) of dimension $n\geq 2,$ and if $N\subset M$ is a compact submanifold of codimension one, then $\mathcal{P}=N\times N$ is an admissible product boundary condition.
@@
(b) and (c). Here the normal exponential map of the submanifold $\partial M$ resp. $N$ is defined $\exp^{\perp}:\nu^{+}(\partial M)\longrightarrow M,$ resp. $\exp^{\perp}:\nu(N)\longrightarrow M.$ The set $\nu^{+}(\partial M)$ is the subset of the normal bundle of $\partial M$ of normal vectors pointing inward and $\nu(N)$ is the normal bundle of $N\subset M.$ Then there is a positive and continuous function $g\in\mathcal{G}^{r}(M)\mapsto\mathrm{inj}^{\perp}(g),$ which is also called the injectivity radius of the normal exponential map $\exp^{\perp},$ such that the restriction onto the normal vectors of length $\mathrm{inj}^{\perp}$ is a diffeomorphism onto its image. Any non-trivial $(g,\partial M\times\partial M)$-geodesic, resp. $(g,N\times N)$-geodesic $\gamma:[0,1]\longrightarrow M,$ satisfies $L_{g}(\gamma)\geq 2\mathrm{inj}^{\perp}(g).$
@@
It is a standard construction that one can add a cylindrical neck to the manifold $M$ with smooth boundary, i.e. there exists a Riemannian manifold $M_{0}$ with metric $g_{0}$ such that $M$ is an open submanifold and $g$ is the restriction of $g_{0}$ to $M.$ Furthermore, for some positive $\epsilon>0$ the normal exponential map of the hypersurface $\partial M$ defines an injective immersion $\phi:(t,x)\in[-\epsilon,0]\times\partial M\longmapsto M_{0}.$ The Riemannian metric on $M$ near $\partial M$ then has the form $dt^{2}+h_{t}$ for a smooth family $t\in[-\epsilon,0]\mapsto h_{t}\in\mathcal{G}(\partial M)$ of Riemannian metrics $h_{t}$ on $\partial M.$ Then we define the manifold $M_{0}$ as the disjoint union of $M$ and the cylinder $\epsilon,1]\times\partial M\ni(t,x)$ where we identify $(t,x)$ and $\phi(t,x)$ for all $(t,x)\in(-\epsilon,0]\times\partial M.$ We extend the Riemannian metric on $\epsilon,1]\times\partial M$ as $dt^{2}+h_{t}$ such that $h_{t}=h_{1}$ for all $t\in[2/3,1].$ Hence the set $\phi([2/3,1]\times\partial M)$ is a cylindrical end, which is isometric to the Riemannian product $2/3,1]\times(\partial M,h_{1}).$ This construction can be found for example in [10, Section 5].
We take two copies of $(M_{0},g_{0})$ and identify them along the boundary $\partial M_{0}.$ This way we obtain a compact Riemannian manifold $(M_{1},g_{1})$ which consists of two copies of $(M_{0},g_{0})$ without boundary, such that $\partial M_{0}$ is a totally geodesic submanifold, which is the fixed point set of an isometric involution interchanging the two copies of $(M_{0},g_{0}).$
@@
In the setting of Lemma 5 let $\phi:\mathcal{G}(M_{1})\longrightarrow\mathcal{G}(M)$ be the map defined by restricting the Riemannian metric $g$ on $M_{1}$ to the submanifold $M,$ i.e. $\phi(g)=g|M.$ The following statements are an immediate consequence of the definition of the strong topology, in particular of the definition of a strong basic neighborhood cf. [14, Section 2.1] applied to the space of Riemannian metrics on $M$ resp. $M_{1}.$ After fixing a Riemannian metric the space $\mathcal{G}(M)$ of Riemannian metrics of class $C^{k},k\geq 2$ on the manifold $M$ endowed with with the strong topology can be seen as an open subset in the Banach space of $C^{k}$-sections of the bundle of symmetric bilinear forms on the manifold $M.$ If $\mathcal{U}$ is an open and dense subset of $\mathcal{G}(M_{1})$ then $\phi(\mathcal{U})$ is an open and dense subset of $\mathcal{G}(M).$ And if $\mathcal{V}$ is a residual subset of $\mathcal{G}(M_{1}),$ then the restriction $\phi(\mathcal{V})$ is a residual subset of $\mathcal{G}(M).$
@@
(b) Let $(M,g)$ a compact Riemannian manifold of dimension $n\geq 2$ and let $N\subset M$ be a compact submanifold of codimension $1.$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics on $M$ such that for any $g\in\mathcal{G}^{*}[L]$ no $(g,N\times N)$-geodesic of length $\leq L$ is part of a closed geodesic.
@@
For $L>0$ let $\mathcal{G}^{*}[L]$ be the set of Riemannian metrics for which all closed geodesics with energy $\leq L$ are non-degenerate and simple, cf. [23, Thm.1] such that no restriction onto a subinterval defines - up to linear reparametrization - a geodesic segment between $p$ and $q$ in case (a) resp. a $(g,N\times N)$-geodesic in case (b). This set is open, we have to show that it is a dense subset of $\mathcal{G}(M).$ A metric $g\in\mathcal{G}^{*}[L]$ has only finitely many geometrically distinct prime closed geodesics $c_{j}:S^{1}\longrightarrow M,j=1\ldots,k$ of energy $\leq L$ and there is an open subset $\mathcal{U}\subset\mathcal{G}^{*}[L]$ and continuous functions $g\in\mathcal{U}\mapsto v_{j}(g)\in TM$ such that the curve $t\in\mathbb{R}\mapsto c_{j}(t)=\Phi_{g}^{t}(v_{j}(g)),t\in S^{1}$ is a closed geodesic of $g$ and there are - up to geometric equivalence - no further closed geodesics of length $\leq L$ of the metric $g,$ cf. [3, p.12-13], resp. the Proof of Part (b) of Theorem 4. Here $\Phi^{t}_{g}:TM\longrightarrow TM,t\in\mathbb{R}$ denotes the geodesic flow, i.e. for $v\in T_{p}M,t\in\mathbb{R}\mapsto\Phi_{g}^{t}(v)\in TM$ is the periodic orbit in the tangent bundle whose projection on the manifold is the geodesic $c:\mathbb{R}\longrightarrow M$ with initial point $p=c(0)$ and initial direction $v=c^{\prime}(0).$
@@
Let $D^{n-1}_{2\eta}=\{x\in\mathbb{R}^{n-1}\,;\,\|x\|\leq 2\eta\}$ be the Euclidean (n1)(n-1)-disc of radius $2\eta$ for $\eta>0.$ On $(t,x)=(t,x_{1},\ldots,x_{n-1})\in[-2\eta,2\eta]\times D^{n-1}_{2\eta}$ a Riemannian metric is given by
@@
$\frac{ds^{2}+2\epsilon f(t)dsdx_{1}+\epsilon^{2}f^{2}(t)dx_{1}^{2}}{(1-\epsilon f^{\prime}(t))^{2}}+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}\,.$
@@
is sufficiently close go $g$ in the $C^{k}$-topology for $k\geq 2$ and the the curve $x_{1}=\ldots=x_{n-1}=0$ coincides with $\gamma$ and is a geodesic of $\overline{g}$ parametrized by arc length. At the point $p=\gamma(0)$ the curve is orthogonal to the hypersurface $s=0$ but not to the hypersurface $\Sigma=\{t=0\}.$
With the help of Lemma 6 we can prove the following main result, from which we can derive the theorems stated in the Introduction:
@@
(c) Let $M$ be a compact differentiable manifold (without boundary) of dimension $n\geq 3$ and let $N\subset M$ be a compact submanifold of codimension $1.$ Then for a $C^{k}$-generic metric on $M$ with $k\geq 2$ all $(g,N\times N)$-geodesics are non-degenerate and simple, and any two geometrically distinct $(g,N\times N)$-geodesics do not intersect.
@@
It follows from Lemma 6 and from the result [4, Thm.5.1], cf. also [4, p.347], that for $L>0$ the set $\mathcal{G}_{\mathcal{P}}(M)[L]$ of metrics $g\in\mathcal{G}_{\mathcal{P}}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and are not part of a closed geodesic is an open and dense subset of $\mathcal{G}_{\mathcal{P}}=\mathcal{G}_{\mathcal{P}}(M).$ This holds since in all three cases (a),(b),(c) the boundary conditions $\mathcal{P}$ are non-degenerate and admissible, cf. Lemma 4.
Here one has to note that in case (b) one has to consider the differentiable manifold $M$ to be a subset of a compact manifold $M_{1}$ without boundary, cf. Lemma 5. Then one can apply [4, Thm.5.1] and conclude that the set $\mathcal{G}_{\partial M\times\partial M}(M_{1})[L]$ of metrics $g\in\mathcal{G}(M_{1})$ for which all $(g,\partial M\times\partial M)$-geodesics of length $\leq L$ are non-degenerate is an open and dense subset of $\mathcal{G}_{\partial M\times\partial M}(M_{1}).$ Then the set $\mathcal{G}_{\partial M\times\partial M}(M)$ contains all restrictions $g|M$ of metrics $g\in\mathcal{G}_{\partial M\times\partial M}(M_{2})[L],$ and is therefore also dense in $\mathcal{G}(M),$ cf. Remark 3.
@@
We conclude from [12, Thm.2.4] that the energy functional $E_{g}:c\in\Omega_{\mathcal{P}}(M)\longmapsto E_{g}(c)\in\mathbb{R}$ satisfies the Palais-Smale condition. We can apply this result since $M$ is compact, hence any Riemannian metric on $M$ is complete.
@@
The self-adjoint operator $H_{c}$ representing the hessian $d^{2}E_{g}(c)$ at a $(g,\mathcal{P})$-geodesic $c$ is of the form $H_{c}=\mathbbm{1}+k_{c},$ i.e. the sum of the identity and a compact operator $k_{c}.$ A proof can be found in[15, Lem.2.5.2] or in [4, Sec.5.1], where one can set $g_{0}=g_{R},$ hence $A=\mathbbm{1}.$ Since 0 is not an eigenvalue the inverse $H_{c}^{-1}$ is a bounded linear operator. Therefore we can use the implicit function theorem for Banach manifolds, cf. [7, Thm.15.1] and conclude: There is a neigbhourhood $\mathcal{U}$ of $g$ and continuous functions $g\in\mathcal{U}\longmapsto c_{j}[g]\in\Omega_{\mathcal{P}}(M),j=1,\ldots,k$ such that $c_{j}[g]$ is a $(g,\mathcal{P})$-geodesic of $g\in\mathcal{U}.$
One can choose the open neighborhhod $\mathcal{U}$ sufficiently small such that there are no further $(g,\mathcal{P})$-geodesics of $g$ with $E_{g}(\gamma)\leq L,$ cf. the argument in [3, p.12-13].
@@
We first assume $p=q.$ For a Riemannian metric $g\in\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for some $L>0$ there are only finitely many geometrically distinct and non-degenerate geodesic loops $\gamma_{1},\ldots,\gamma_{N}$ of length $\leq L.$ Assume that - after a possible renumbering - the geodesic loops $\gamma_{1},\ldots,\gamma_{m},m\leq N$ are the prime geodesic loops of length $\leq L.$ Here a geodesic loop $\gamma:[0,1]/\{0,1\}\longrightarrow M$ is prime, if there is no $t_{1}\in(0,1)$ with $p=c(t_{1}).$ Let $v_{j,l}=\gamma_{j}^{\prime}(l),j=1,\ldots,m,l\in\{0,1\}$ be the sequence of tangent vectors of the prime geodesic loops at $p.$ We use the perturbation result Lemma 3 which shows that there is an open and dense subset $\mathcal{G}_{p}[L]\subset\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for which all geodesic loops at $p$ of length $\leq L$ are prime. Note here that if $m<N,$ then the geodesic loops $\gamma_{m+1},\ldots,\gamma_{N}$ are not prime, but their prime parts are loops $\gamma_{j},1\leq j\leq m.$ Since after the perturbation these prime geodesic loops cannot be concatenated to produce a non-prime geodesic loop the perturbed loops $\gamma_{j}[g],j\geq m+1$ have to be prime, too.